Eriaz:

I need to know the x and y intercept of this equation: y=5x-1

Luigi0210:

To find the x intercept, set y equal to 0: \[\large 0=5x-1 \] Solve: \[\large 1=5x \] \[\large \frac{1}{5}=x \] So your x intercept is \(\large \color{red}{(\frac{1}{5}, 0)} \) To find the y intercept, set x=0: \[\large y=5(0)-1 \] giving you: \[\large y=-1\] So your x intercept would be \(\large \color{yellow}{(0, -1)}\)